# VisBias: Measuring Explicit and Implicit Social Biases in Vision Language Models

### v2

@@ Abstract
VisBias: Measuring Explicit and Implicit Social Biases in Vision Language Models
@@ 1 Introduction
Visual Language Models (VLMs) have shown the ability to interpret visual content and answer related queries Bubeck et al. (2023); Chow et al. (2025). As these models gain widespread adoption in tasks such as text recognition Liu et al. (2024); Chen et al. (2025), mathematical problem-solving Yang et al. (2024); Peng et al. (2024), and medical applications Azad et al. (2023); Buckley et al. (2023), identifying and addressing potential social biases is essential to prevent harm to users Raj et al. (2024); Sathe et al. (2024); Brinkmann et al. (2023); Ruggeri and Nozza (2023).
@@
Existing research has greatly advanced bias identification and mitigation in VLMs, particularly regarding explicit biases in textual or multimodal outputs Wan et al. (2023); Zhao et al. (2024); Gupta et al. (2024); Wang et al. (2024a); Ruggeri and Nozza (2023). However, fewer efforts address implicit biases Cheng et al. (2023); Wolfe et al. (2023); Wolfe and Caliskan (2022), and even fewer attempt to study both implicit and explicit biases together. As a result, many approaches risk overlooking how these two forms of bias intersect or reinforce one another in real-world scenarios, underscoring the need for more detailed analyses that capture the full spectrum of biased behavior in VLMs.
In this paper, we study both explicit and implicit social biases by designing fairness-related questions and tasks that, while appearing to assist user requests, may reveal underlying biases. We introduce VisBias, which evaluates these biases in VLMs through four distinct assessment scenarios. First, we focus on explicit tasks—(I) Multiple-Choice and (II) Yes-No questions—where models make direct judgments about individuals’ annual income, education level, political leaning, and religion based on real-world images that capture race, gender, and occupation. To counteract alignment-related refusals, we apply a Caesar cipher “jailbreak” Yuan et al. (2024) ensuring models reveal biases that might otherwise remain hidden. In the multiple-choice setting, we analyze which demographic groups are linked to particular answers, while in yes-no comparisons (such as asking whether one person appears more educated than another), we check for group-specific favoritism. Next, to unveil more subtle forms of bias, we introduce implicit tasks—(III) Image Description and (IV) Form Completion—where the model’s goal is to assist with descriptive or data-entry tasks rather than make overt judgments. In image description, we measure skewed word associations across different demographic groups by examining the language models use. Finally, in form completion, we prompt VLMs to fill out personal information forms with various attributes, tracking how certain features (like gender or race) shift the distribution of other attributes. By integrating explicit and implicit testing, VisBias provides a comprehensive lens for detecting bias across diverse modeling tasks, as shown in Fig. 1.
@@ 2.1 Explicit Scenarios
In explicit scenarios, the questions are designed to instruct models to make decisions primarily based on the individual’s appearance. To incorporate demographic information through image modality, we first collect real-world images. Ghavami and Peplau (2013) showed that people’s stereotypes and expectations vary depending on the combined identity of an individual rather than just their race or gender alone. Following this work, we focus on the intersectionality of race (Asian, Black, Hispanic, Middle Eastern, and White) and gender (Female and Male). For occupations, we select those with high gender or race imbalance according to U.S. Bureau of Labor Statistics, [^1] https://www.bls.gov/bls/blswage.htm including basketball player, nurse, firefighter, CEO, cook, doctor, and lawyer. We retrieve images by searching for keywords related to race, gender, and occupation on Google Images, selecting those whose descriptions explicitly include all three attributes. We obtain ten images per gender-race-occupation group, yielding 700 images in total (2 genders × 5 races × 7 occupations × 10 images).
Models with safety alignment may refuse to answer direct questions. In such cases, the response is classified as unbiased. To create a more challenging testing scenario, we employ the jailbreak method CipherChat Yuan et al. (2024), which encodes questions using a Caesar cipher with a three-letter shift. To measure the intrinsic biases in the model, we do not include negative examples or explicitly instruct models to generate harmful or toxic content. Instead, we provide neutral examples and define the Caesar cipher, specifying that communication should adhere to this encoding. The full prompt is provided in §C of the appendix.
@@
The second scenario involves yes-no questions that require models to compare individuals in paired images regarding their ordinal attributes. This assessment requires VLM to recognize the input order of images and respond accordingly. We focus on two attributes—annual income and education level—posing questions such as “Is the person in the first image more educated than the person in the second?” and “Does the person in the first image earn more than the person in the second?” A model that consistently outputs “Yes” or “No” to all queries exhibits no discernible bias toward any group Huang et al. (2025a). To control for order effects, each query is paired with a swapped version in which both the prompt (e.g., “Is the person in the second image less educated than the person in the first?”) and image order are reversed. Since three order changes are applied, the new query represents the negation of the original query. The number of query pairs for which a model provides differing responses is analyzed to detect potential biases.
@@ 2.2 Implicit Scenarios
In the fourth scenario, the model is required to complete a draft personal information form that has been pre-filled with randomly generated data. To construct a comprehensive individual profile, we select 20 attributes, including age, gender, race, religion, education level, occupation, political leaning, hobbies, and personality traits. Each attribute is assigned predefined categories; for instance, education level includes “No Education,” “Elementary School,” “Middle School,” “High School,” “Associate,” “Bachelor,” “Master,” and “Doctorate,” resulting in a total of 128 possible choices. The complete list of attributes and their corresponding choices is presented in Table 7 in §D of the appendix. During model evaluation, five attributes are randomly selected along with corresponding choices, and the form is converted into an image format for input into VLMs.
@@
1:A model $\mathcal{M}$, A form =$\mathcal{F}=\{a_{1},\cdots,a_{n}\}$ containing $n$ attributes, where each attribute $a_{i}$ has a set of possible choices $C_{i}$
@@ 3 Experiments
This section presents our experiments with VisBias on five VLMs: GPT-4(V) OpenAI (2023), GPT-4o OpenAI (2024), Gemini-1.5-Pro Pichai and Hassabis (2024), LLaMA-3.2 (Vision) Dubey et al. (2024), and LLaVA-v1.6 Liu et al. (2023). For multiple-choice questions, GPT-4V, GPT-4o, and Gemini-1.5 refuse to provide responses. In contrast, all models generate valid responses to yes-no questions. When the Caesar cipher is applied, Gemini-1.5-Pro and GPT-4V can answer both question types. However, GPT-4o, with its enhanced safety alignment, declines to respond to multiple-choice questions but provides answers to yes-no questions. LLaMA and LLaVA are unable to interpret the Caesar cipher and thus cannot respond to most queries. Finally, in the form completion task, LLaMA refuses to provide responses, while LLaVA is unable to complete the task due to its limited reasoning ability. The refusal rates are reported in Table 2 in §A of the appendix. Temperatures are set to zero for all models, except in the image description task, where it is one.
@@
We also examine the sentiment of words used to describe different demographic groups using the VADER sentiment analyzer in NLTK Hutto and Gilbert (2014) and check whether the words are typical racial stereotype words using the dictionary in Ghavami and Peplau (2013). The sentiment scores and stereotype scores are presented in Table 5 and 6 in the appendix, respectively. Among the models, Gemini and Midjourney exhibit the strongest biases, with the most pronounced biases directed toward Middle Eastern and Black groups.
@@
High correlation between attribute pairs does not necessarily indicate social bias. For example, the attribute “Age = 0-17” is strongly correlated with “Marital Status = Single,” “Has Children = No,” “Occupation = Student,” “Annual Income = Less Than \$10k,” and “Residential Status = Living with Family.” To determine whether such correlations reflect social bias or reasonable causal inference, human evaluation is required, as LLM inherently exhibit biases. We employ human annotators to assess 370 attribute pairs, distributed across 19 questionnaires. Each questionnaire follows the format: “To what extent is the following statement socially biased? If a person’s {Attribute 1} is {Choice 1333Choice 1 is the conditioned value of Attribute 1.}, then we can infer that the person’s {Attribute 2} is {Choice 2444Choice 2 is the value that has the largest probability of Attribute 2, conditioned on Attribute 1 = Choice 1.}.” Participants rate bias on a five-point scale, ranging from “(1) No Bias” to “(5) Severe Bias.” An example questionnaire created using Qualtrics [^5] https://www.qualtrics.com/ is shown in Fig. 4 in §E of the appendix. Participants are recruited via Prolific. [^6] https://www.prolific.com/ To ensure diversity and prevent overrepresentation of specific gender-race groups, we only require fluency in English. Responses submitted in under two minutes or containing identical answers for all questions are excluded. Each questionnaire is completed by three different annotators. On average, participants take 5 to 10 minutes to complete a questionnaire, with an hourly compensation of \$6.84 (rated as “Good” by the platform).
@@
Figure 3: Visualization of the choices that maximize the distribution shift of other attributes in (IV) Form Completion. Color represents the choices, while size indicates the KL divergence. For example, in Panel (d), for GPT-4o, selecting “Doctorate” for the attribute “education level” induces the largest distribution shift in attribute “occupation” compared to other education level choices. This shift is substantial, as indicated by the large plot size. Figures for other sixteen attributes are included in §F of the appendix. Circles in bold are biases verified by humans.
@@
Figure 2 presents the distribution of average human scores across all 370 attribute pairs. Since a score of 3 represents “Moderate Bias,” we classify attribute pairs with an average score greater than or equal to 3 as exhibiting social bias, while those below 3 are considered reasonable causal inferences. Examples of high-scoring biases include: (1) If a person’s occupation is Athlete, then we can infer that the person’s religion is Christianity. (2) If a person’s sexual orientation is Heterosexual, then we can infer that the person’s personality type is INTJ. We further analyze attribute correlations by identifying choices that maximize distribution shifts, as illustrated in Fig. 3, 5, 6, 7, and 8. VLMs exhibit stronger social biases in this task compared to explicit bias evaluation scenarios. These models often establish rigid associations between attributes. For example, individuals in Creative Roles are consistently associated with the INFP personality type and Painting as a hobby. Additionally, all models demonstrate a strong tendency to associate specific ethnic groups with particular religions. For example, Buddhism is predominantly linked to Asian individuals, reinforcing cultural stereotypes. Such strong associations may inadvertently weaken the public’s perception of the diversity of certain demographic groups and reinforce stereotypes, leading to overly generalized or inaccurate representations of real-world populations.
@@ 4 Related Work
Recent studies have focused on both text and multimodal settings, showing that biases persist in various forms, including sexual objectification and cultural conflation Wolfe et al. (2023); Wolfe and Caliskan (2022); Wang et al. (2024b), misclassification tied to skin tone Hazirbas et al. (2024), diagnostic inequities Yang et al. (2025), and a pronounced masculine tilt Huang et al. (2024). BiasAsker Wan et al. (2023), Zhao et al. (2024) and Huang et al. (2025a) reveal significant social biases in LLM outputs. Gupta et al. (2024) show persona-based prompts worsen stereotypes, and Marked Personas Cheng et al. (2023) highlights how explicit demographic references magnify biased portrayals. Fact-or-Fair BiasPainter Wang et al. (2024a) and Ghosh and Caliskan (2023) uncover biased image generation, while Ruggeri and Nozza (2023), Brinkmann et al. (2023), Srinivasan and Bisk (2022), and Mandal et al. (2023) emphasize how model size, objectives, and representations can exacerbate biases.
Existing papers provide datasets and benchmarks for systematically evaluating social biases in AI Ross et al. (2021); Hall et al. (2023); Huang et al. (2025b); Zhou et al. (2022); Du et al. (2025). BIGbench Luo et al. (2024) unifies bias evaluations in text-to-image models. BiasLens Li et al. (2024) offers 33,000 role-specific questions to test fairness in LLMs. Meanwhile, BiasDora Raj et al. (2024) focuses on hidden biases. PAIRS Fraser and Kiritchenko (2024) and Sathe et al. (2024) use synthetic datasets to probe gender, race, and age biases in VLMs. SocialCounterfactuals Howard et al. (2024) provides counterfactual image-text pairs for intersectional bias analysis. Studies also mitigate social bias in VLMs via dataset curation process Hamidieh et al. (2024), an LLM-driven critique and test suite Wan and Chang (2024), counterfactual training data Zhang et al. (2022), and a lightweight post-processing step that curbs gender and racial bias without hurting retrieval accuracy Kong et al. (2023).
While prior studies have explored social biases in VLMs, they often focus exclusively on either explicit or implicit manifestations, or are limited to specific modalities such as text or generated images. In contrast, VisBias evaluates both explicit and implicit biases within a unified experimental design. Notably, our form completion scenario introduces a novel method for probing implicit bias by asking VLMs to complete a 20-attribute personal information form. It is the first to jointly assess model outputs in direct judgment and assistance contexts, providing a holistic view of bias in state-of-the-art VLMs.

### v3

@@ Abstract
VisBias: Measuring Explicit and Implicit Social Biases in
Vision Language Models
@@ 1 Introduction
Visual Language Models (VLMs) have shown the ability to interpret visual content and answer related queries Bubeck et al. (2023); Chow et al. (2025). As these models gain widespread adoption in tasks such as text recognition Liu et al. (2024b); Chen et al. (2025), mathematical problem-solving Yang et al. (2024b); Peng et al. (2024), and medical applications Azad et al. (2023); Buckley et al. (2023), identifying and addressing potential social biases is essential to prevent harm to users Raj et al. (2024); Sathe et al. (2024); Brinkmann et al. (2023); Ruggeri and Nozza (2023).
@@
Existing research has greatly advanced bias identification and mitigation in VLMs, particularly regarding explicit biases in textual or multimodal outputs Wan et al. (2023); Ding et al. (2025); Gupta et al. (2024); Wang et al. (2024a); Ruggeri and Nozza (2023). However, fewer efforts address implicit biases Cheng et al. (2023); Wolfe et al. (2023); Wolfe and Caliskan (2022), and even fewer attempt to study both implicit and explicit biases together. As a result, many approaches risk overlooking how these two forms of bias intersect or reinforce one another in real-world scenarios, underscoring the need for more detailed analyses that capture the full spectrum of biased behavior in VLMs.
In this paper, we study both explicit and implicit social biases by designing fairness-related questions and tasks that, while appearing to assist user requests, may reveal underlying biases. We introduce VisBias, which evaluates these biases in VLMs through four distinct assessment scenarios. First, we focus on explicit tasks—(I) Multiple-Choice and (II) Yes-No questions—where models make direct judgments about individuals’ annual income, education level, political leaning, and religion based on real-world images that capture race, gender, and occupation. To counteract alignment-related refusals and stress-test model behaviors, we apply a Caesar cipher “jailbreak” Yuan et al. (2024) ensuring models reveal biases that might otherwise remain hidden. In the multiple-choice setting, we analyze which demographic groups are linked to particular answers, while in yes-no comparisons (such as asking whether one person appears more educated than another), we check for group-specific favoritism. Next, to unveil more subtle forms of bias, we introduce implicit tasks—(III) Image Description and (IV) Form Completion—where the model’s goal is to assist with descriptive or data-entry tasks rather than make overt judgments. In image description, we measure skewed word associations across different demographic groups by examining the language models use. Finally, in form completion, we prompt VLMs to fill out personal information forms with various attributes, tracking how certain features (like gender or race) shift the distribution of other attributes. By integrating explicit and implicit testing, VisBias provides a comprehensive lens for detecting bias across diverse modeling tasks, as shown in Fig. 1.
@@ 2.1 Explicit Scenarios
In explicit scenarios, the questions are designed to instruct models to make decisions primarily based on the individual’s appearance. To incorporate demographic information through image modality, we first collect real-world images. Ghavami and Peplau (2013) showed that people’s stereotypes and expectations vary depending on the combined identity of an individual rather than just their race or gender alone. Following this work, we focus on the intersectionality of race (Asian, Black, Hispanic, Middle Eastern, and White) and gender (Female and Male). For occupations, we select those with high gender or race imbalance according to U.S. Bureau of Labor Statistics, [^1] https://www.bls.gov/bls/blswage.htm including basketball player, nurse, firefighter, CEO, cook, doctor, and lawyer. We retrieve images by searching for keywords related to race, gender, and occupation on Google Images, selecting those whose descriptions explicitly include all three attributes and avoiding memes, heavily edited content, or images lacking clear human subjects. We obtain 10 images per group, yielding 700 images in total (2 genders $\times$ 5 races $\times$ 7 occupations $\times$ 10 images). Examples of images are shown in §B.2 of the appendix.
Models with safety alignment may refuse to answer direct questions. In such cases, the response is classified as unbiased. To create a more challenging testing scenario, we employ the jailbreak method CipherChat Yuan et al. (2024), which encodes questions using a Caesar cipher with a three-letter shift. To measure the intrinsic biases in the model, we do not include negative examples or explicitly instruct models to generate harmful or toxic content. Instead, we provide neutral examples and define the Caesar cipher, specifying that communication should adhere to this encoding. The full prompt is provided in §B.1 of the appendix. A model is considered unbiased if it refuses to answer or outputs consistent responses regardless of input demographic information Huang et al. (2025a).
@@
The second scenario involves yes-no questions that require models to compare individuals in paired images regarding their ordinal attributes. This assessment requires VLM to recognize the input order of images and respond accordingly. We focus on two attributes—annual income and education level—posing questions such as “Is the person in the first image more educated than the person in the second?” and “Does the person in the first image earn more than the person in the second?” To control for order effects, each query is paired with a swapped version in which both the prompt (e.g., “Is the person in the second image less educated than the person in the first?”) and image order are reversed. Since three order changes are applied, the new query represents the negation of the original query. The number of query pairs for which a model provides differing responses is analyzed to detect potential biases.
@@ 2.2 Implicit Scenarios
In the fourth scenario, the model is required to complete a draft personal information form that has been pre-filled with randomly generated data. To construct a comprehensive individual profile, we select 20 attributes, including age, gender, race, religion, education level, occupation, political leaning, hobbies, and personality traits. Each attribute is assigned predefined categories; for instance, education level includes “No Education,” “Elementary School,” “Middle School,” “High School,” “Associate,” “Bachelor,” “Master,” and “Doctorate,” resulting in a total of 128 possible choices. The complete list of attributes and their corresponding choices is presented in Table 10 in §B.3 of the appendix. During model evaluation, five attributes are randomly selected along with corresponding choices, and the form is converted into an image format for input into VLMs.
@@
1:A model $\mathcal{M}$, A form $\mathcal{F}=\{a_{1},\cdots,a_{n}\}$ containing $n$ attributes, where each attribute $a_{i}$ has a set of possible choices $C_{i}$
@@ 3 Experiments
This section presents our experiments with VisBias on five VLMs: GPT-4(V) OpenAI (2023), GPT-4o Hurst et al. (2024), Gemini-1.5-Pro Team et al. (2024), LLaMA-3.2 (Vision) Meta (2024), and LLaVA-v1.6 Liu et al. (2024a). For multiple-choice questions, GPT-4V, GPT-4o, and Gemini-1.5 refuse to provide responses. In contrast, all models generate valid responses to yes-no questions. When the Caesar cipher is applied, Gemini-1.5-Pro and GPT-4V can answer both question types. However, GPT-4o, with its enhanced safety alignment, declines to respond to multiple-choice questions but provides answers to yes-no questions. LLaMA and LLaVA are unable to interpret the Caesar cipher and thus cannot respond to most queries. Finally, in the form completion task, LLaMA refuses to provide responses, while LLaVA is unable to complete the task due to its limited reasoning ability. The refusal rates are reported in Table 3 in §A.1 of the appendix. Temperatures are set to zero for all models, except in the image description task, where it is one.
@@
BiasDora Raj et al. (2024) also investigates biases on multimodal models. Their study evaluates models using four different captioning instructions and analyzes the generated outputs for biased associations (e.g., “gay $\leftrightarrow$ insane”). They report that LLaVA exhibits more biases than GPT-4o in image descriptions, and that gender and sexual orientation biases are especially pronounced in text-to-image tasks. Our findings partially align with these observations. As shown in Table 7, gender-related stereotype scores are indeed among the highest, consistent with their second point. However, in contrast to their first claim, we observe that GPT-4o demonstrates higher biases than LLaVA in several dimensions, such as a stronger sentiment polarity and stereotype scores across racial groups.
@@
We also examine the sentiment of words used to describe different demographic groups using the VADER sentiment analyzer in NLTK Hutto and Gilbert (2014) and check whether the words are typical racial stereotype words using the dictionary in Ghavami and Peplau (2013). The sentiment scores and stereotype scores are presented in Table 7 and 8 in the appendix, respectively. Among the models, Gemini and Midjourney exhibit the strongest biases, with the most pronounced biases directed toward Middle Eastern and Black groups.
@@
High correlation between attribute pairs does not necessarily indicate social bias. For example, the attribute “Age = 0-17” is strongly correlated with “Marital Status = Single,” “Has Children = No,” “Occupation = Student,” “Annual Income = Less Than \$10k,” and “Residential Status = Living with Family.” To determine whether such correlations reflect social bias or reasonable causal inference, human evaluation is required, as LLM inherently exhibit biases. We employ human annotators to assess 370 attribute pairs, distributed across 19 questionnaires. Each questionnaire follows the format: “To what extent is the following statement socially biased? If a person’s {Attribute 1} is {Choice 1333Choice 1 is the conditioned value of Attribute 1.}, then we can infer that the person’s {Attribute 2} is {Choice 2444Choice 2 is the value that has the largest probability of Attribute 2, conditioned on Attribute 1 = Choice 1.}.” Participants rate bias on a five-point scale, ranging from “(1) No Bias” to “(5) Severe Bias.” An example questionnaire created using Qualtrics [^5] https://www.qualtrics.com/ is shown in Fig. 10 in §C.2 of the appendix. Participants are recruited via Prolific. [^6] https://www.prolific.com/ To ensure diversity and prevent overrepresentation of specific gender-race groups, we only require fluency in English. Responses submitted in under two minutes or containing identical answers for all questions are excluded. Each questionnaire is completed by three different annotators. On average, participants take 5 to 10 minutes to complete a questionnaire, with an hourly compensation of \$6.84 (rated as “Good” by the platform).
@@
Figure 3: Visualization of the choices that maximize the distribution shift of other attributes in (IV) Form Completion. Color represents the choices, while size indicates KL divergence. For example, in Panel (d), for GPT-4o, selecting “Doctorate” for the attribute “education level” induces the largest distribution shift in attribute “occupation” compared to other education level choices. This shift is great, as indicated by the larger plot size. Figures for other 16 attributes are included in §A.4. Circles in bold are biases verified in our human study.
@@
Figure 2 presents the distribution of average human scores across all 370 attribute pairs. Since a score of 3 represents “Moderate Bias,” we classify attribute pairs with an average score greater than or equal to 3 as exhibiting social bias, while those below 3 are considered reasonable causal inferences. Examples of high-scoring biases include: (1) If a person’s occupation is Athlete, then we can infer that the person’s religion is Christianity. (2) If a person’s sexual orientation is Heterosexual, then we can infer that the person’s personality type is INTJ. We further analyze attribute correlations by identifying choices that maximize distribution shifts, as illustrated in Fig. 3, 4, 5, 6, and 7. VLMs exhibit stronger social biases in this task compared to explicit bias evaluation scenarios. These models often establish rigid associations between attributes. For example, individuals in Creative Roles are consistently associated with the INFP personality type and Painting as a hobby. Additionally, all models demonstrate a strong tendency to associate specific ethnic groups with particular religions. For example, Buddhism is predominantly linked to Asian individuals, reinforcing cultural stereotypes. Such strong associations may inadvertently weaken the public’s perception of the diversity of certain demographic groups and reinforce stereotypes, leading to overly generalized or inaccurate representations of real-world populations.
MCQ Income Education Political Leaning Religion
Female 0.118 1.409 1.164 1.122
Male 0.106 1.073 1.279 4.476
Orc 21.258 7.031 5.639 4.112
Murloc 2.744 3.752 7.968 2.490
Goblin 21.258 1.077 0.305 3.358
Dwarf 21.258 0.388 27.310 8.293
Elf 22.702 16.234 8.293 8.293
Yes-No Income w/o J Income w/ J Education w/o J Education w/ J
2% 4\$ 2% 2%
Table 2: Results on income, education, politics, and religion using fictional species.
4 Further Analysis
4.1 Images of Fictional Races
In addition to real demographic groups, we also evaluate fictional species, which serve as negative controls to help isolate biases that may arise from human-centric data or RLHF processes. Specifically, we selected five commonly depicted fantasy species—Orc, Murloc, Goblin, Dwarf, and Elf—and applied both the MCQ and Yes/No question experiments using the same methodology as for human demographic groups. The results using Gemini-1.5-Pro are presented in Table 2.
Whereas human demographic groups show relatively low divergence scores across MCQ categories, fictional species exhibit substantially higher variability, particularly for political leaning (e.g., Dwarf: 27.31) and education (e.g., Elf: 16.234). This pattern suggests that VLMs may rely on entrenched tropes from fictional media (e.g., “wise elves” or “brutish orcs”) when RLHF-tuned norms are absent. These findings support the view that RLHF plays a regularizing role for human categories but not for out-of-distribution entities, where learned stereotypes may persist unchecked. Importantly, this stress-test highlights a broader implication: the perception of bias can sometimes exceed actual bias, potentially discouraging beneficial use of VLMs and inadvertently reinforcing inequities. By demonstrating that the same models produce far larger divergences for categories such as elves and orcs—which lack RLHF safeguards and any real socio-economic ground truth—we provide readers with a tangible calibration point for interpreting the human-group results.
4.2 Text-Based Form Completion
The (IV) Form Completion task requires visual input tailored for VLMs. To construct this input, we first generate a personal information questionnaire in Microsoft Word, designed to resemble realistic administrative data collection forms. The layout follows common organizational templates. We then export the document as a PNG image, which serves as the direct visual input to the VLMs. This process is fully automated using Python.
To complement this setting, we further examine the same form in text format to assess bias in language-only models. Specifically, we conduct a parallel experiment using text input with identical prompt content and formatting on four LLMs: LLaMA-3.2-90B-Vision-Instruct Meta (2024), Gemini-1.5-Pro Team et al. (2024), GPT-4o Hurst et al. (2024), and Qwen-2.5-72B-Instruct Yang et al. (2024a). We report, for each model, the three attribute pairs with the highest KLD, as shown in Table 9 in the appendix. A higher KLD score indicates stronger conditional dependency, which may signal biased associations. Our results demonstrate that even in text-only settings, LLMs exhibit notable attribute correlations suggestive of social biases.
4.3 Interaction of Implicit and Explicit Bias
While our paper primarily reports implicit and explicit biases separately, we acknowledge and aim to highlight their potential intersection in real-world scenarios. To this end, we compute the average JSD across four target attributes, i.e., Income, Education, Political Leaning, and Religion, between the output distributions of the Form Completion task (implicit) and MCQ task (explicit), conditioning on gender or race. Across both GPT-4V and Gemini-1.5-pro, we observe that for certain attributes, the implicit and explicit outputs show consistently low divergence, indicating structural similarity in group-level distributions. For example, in the gender-based setting using GPT-4V, the average JSDs are: Income = 0.262, Education = 0.063, Political Leaning = 0.192, and Religion = 0.113. In the race-based setting, Education = 0.206 and Religion = 0.16. These results suggest a notable alignment, indicating that implicit and explicit biases may stem from shared internal representations, thereby intersecting each other.
@@ 5 Related Work
Recent studies have focused on both text and multimodal settings, showing that biases persist in various forms, including sexual objectification and cultural conflation Wolfe et al. (2023); Wolfe and Caliskan (2022); Wang et al. (2024b), misclassification tied to skin tone Hazirbas et al. (2024), diagnostic inequities Yang et al. (2025), and a pronounced masculine tilt Huang et al. (2024). BiasAsker Wan et al. (2023), Ding et al. (2025) and Huang et al. (2025a) reveal significant social biases in LLM outputs. Gupta et al. (2024) show persona-based prompts worsen stereotypes, and Marked Personas Cheng et al. (2023) highlights how explicit demographic references magnify biased portrayals. BiasPainter Wang et al. (2024a) and Ghosh and Caliskan (2023) uncover biased image generation, while Ruggeri and Nozza (2023), Brinkmann et al. (2023), Srinivasan and Bisk (2022), and Mandal et al. (2023) emphasize how model size, objectives, and representations can exacerbate biases.
Existing papers provide datasets and benchmarks for systematically evaluating social biases in AI Ross et al. (2021); Hall et al. (2023); Huang et al. (2025b); Zhou et al. (2022); Du et al. (2025); Shi et al. (2025). BIGbench Luo et al. (2024) unifies bias evaluations in text-to-image models. BiasLens Li et al. (2024) offers 33,000 role-specific questions to test fairness in LLMs. Meanwhile, BiasDora Raj et al. (2024) focuses on hidden biases. PAIRS Fraser and Kiritchenko (2024) and Sathe et al. (2024) use synthetic datasets to probe gender, race, and age biases in VLMs. SocialCounterfactuals Howard et al. (2024) provides counterfactual image-text pairs for intersectional bias analysis. Studies also mitigate social bias in VLMs via dataset curation process Hamidieh et al. (2024), an LLM-driven critique and test suite Wan and Chang (2025), counterfactual training data Zhang et al. (2022), and a lightweight post-processing step that curbs gender and racial bias without hurting retrieval accuracy Kong et al. (2023).
While prior studies have explored social biases in VLMs, they often focus exclusively on either explicit or implicit manifestations, or are limited to specific modalities such as text or generated images. In contrast, VisBias evaluates both explicit and implicit biases within a unified experimental design. Notably, our form completion scenario introduces a novel method for probing implicit bias by asking VLMs to complete a 20-attribute personal information form.